# VLBiMan: Vision-Language Anchored One-Shot Demonstration Enables Generalizable Bimanual Robotic Manipulation

### v2

@@ 2 Related Works
Generalizable Representations for Manipulation. Traditional robotic manipulation often relied on structured representations built upon strong priors Kaelbling & Lozano-Pérez (2013); Dantam et al. (2018); Migimatsu & Bohg (2020); Tyree et al. (2022), such as object geometry or rigid-body assumptions, typically via estimating 6D poses or manually specifying grasp configurations. They are hard to scale in unstructured environments. With the rise of data-driven techniques, more flexible representations have emerged, including keypoints Papagiannis et al. (2024); Gao et al. (2024); Wen et al. (2024b); Grannen et al. (2021), affordances Ju et al. (2024); Nasiriany et al. (2024); Zhao et al. (2023b), dynamic flow fields Colomé & Torras (2018); Weng et al. (2022), and invariant object-centric correspondences Ko et al. (2024); Zhang & Boularias (2024); Zhang et al. (2023). Some works further leverage human demonstrations to retarget 3D hand trajectories to robots Chen et al. (2024a); Li et al. (2024); Kerr et al. (2024); Chen et al. (2024b). However, these approaches often depend on private datasets, retraining, or complex retargeting pipelines, limiting scalability. In contrast, our method essentially anchors adaptation to object representive points without retraining, achieving greater efficiency and generality.
@@
Efficient Bimanual Robotic Manipulation. Recent advances in bimanual manipulation have showcased the power of large Vision-Language-Action (VLA) models Black et al. (2024); Liu et al. (2025a); Pertsch et al. (2025) trained on extensive teleoperated demonstrations Fang et al. (2024a); Khazatsky et al. (2024); O’Neill et al. (2024); Bu et al. (2025). However, these approaches are highly suspected of lacking efficiency, as scaling to unseen objects or tasks often requires re-collecting and retraining. Alternative efforts explore leveraging large-scale Internet Ponimatkin et al. (2025); Ye et al. (2025); Bharadhwaj et al. (2024) or egocentric human-hand videos Zhan et al. (2024); Liu et al. (2024b); Grauman et al. (2024); Zhao et al. (2025); Kareer et al. (2024), yet the embodiment gap between human and robot limits direct usability. Some methods improve sample efficiency by learning visuomotor policies Chi et al. (2023); Ze et al. (2024) from a small set of real-world robot data, but their generalization remains limited. While one-shot imitation learning Wen et al. (2022); Bahety et al. (2024); Zhou et al. (2025); Wang & Johns (2025); Mao et al. (2023); Liu et al. (2025b); Biza et al. (2023) reduces data demands, the high-dimensional action space and coordination complexity of bimanual control hinder learning efficiency. In contrast, VLBiMan achieves efficient adaptation from a single bimanual demonstration by leveraging VLMs to handle novel variations, while reusing decomposed task-invariant atomic skills. These lead to both data and computational efficiency.
@@
Large Foundation Models for Robotics. Integrating LLMs and VLMs into robotics is a prominent trend to enable generalizable agents Ma et al. (2025); Huang et al. (2025); Fang et al. (2025); Feng et al. (2025). LLMs are utilized for high-level task understanding and planning, such as decomposing instructions into executable subtasks or generating scripts Liang et al. (2023); Singh et al. (2023); Szot et al. (2024); Huang et al. (2024a). Meanwhile, VLMs facilitate visually grounded perception through semantic prompts, enabling object-level detection and segmentation. For fine-grained tasks, Visual Foundation Models (VFMs) Oquab et al. (2024); Ravi et al. (2025) are further employed to find keypoints Papagiannis et al. (2024); Gao et al. (2024); Wen et al. (2024b) or dense correspondences Ko et al. (2024); Zhang & Boularias (2024). Recent efforts like ReKep Huang et al. (2024b), MOKA Fang et al. (2024b), RobotPoint Yuan et al. (2024a), and RAM Kuang et al. (2024) combine LLMs and VLMs into modular pipelines that follow the perceive-understand-plan-act paradigm to achieve zero-shot generalization. These approaches often rely on engineered prompts and ambiguous intermediate representations (e.g., region of interest or keypoint clusters) requiring additional post-processing. In contrast, VLBiMan avoids LLM-based instruction parsing and task decomposition, which are brittle and labor-intensive. Instead, we build on one-shot demonstrations with precise action labels, using VLMs to extract semantically grounded action structures that are adaptively composed and reused, enabling efficient and scalable bimanual manipulation.
@@ 3.2 Vision-Language Anchored Adaptation
VFM-Based Geometric Feasibility. To adapt grasping or alignment poses, we introduce a three-step process. (1) Firstly, we compute relative position transformation $\Delta\mathbf{T}$ between new object placement and reference demonstration via task-specific representative points (e.g. the 2D mask centroid or a task-specific contact point on the table-facing boundary). Examples of two kinds of representative points can be found in Fig. 3. Let $\mathbf{p}^{\texttt{demo}}$, $\mathbf{p}^{\texttt{new}}$ denote the representative 3D positions back-projected from 2D points via stereo and calibrated camera intrinsics. The relative position shift is $\Delta\mathbf{x}=\mathbf{p}^{\texttt{new}}-\mathbf{p}^{\texttt{demo}}$. (2) To account for orientation-sensitive objects (such as pen, spoon and lying down bottle), we compute the principal axis from second-order image moments Chaumette (2004); Kotoulas & Andreadis (2007) of the 2D mask and derive relative rotation $\Delta\theta=\angle(\mathbf{v}^{\texttt{new}},\mathbf{v}^{\texttt{demo}})$ via angular deviation. The final adapted grasp pose $\widetilde{\mathbf{T}}$ is obtained by applying $(\Delta\mathbf{x},\Delta\theta)$ to the original grasp pose in robot coordinates via calibrated hand-eye transformation. (3) For category-level variation, we measure shape-induced feasibility change through height and width differences. For example, the $z$-extent of the object point cloud $\mathcal{P}_{k}^{\texttt{3D}}$ yields $\Delta h_{k}=\max_{z}(\mathcal{P}_{k}^{\texttt{3D}})-\min_{z}(\mathcal{P}_{k}^{\texttt{3D}})$. This is used to adjust vertical placement motions or inter-arm distances for tools or containers.
@@ 4 Experiments
Baselines and Metric: For each task setting, we conduct 20 trials, where objects are randomly located or replaced, and the success rate will be reported. For baselines, we compare to Robot-ABC Ju et al. (2024) based on keypoint affordance prediction with using AnyGrasp Fang et al. (2023) for initial grasping (After which, the remaining trajectory is obtained by trivial modules combination), as well as ReKep Huang et al. (2024b) based on VFMs (SAM Kirillov et al. (2023) and DINOv2 Oquab et al. (2024)) and GPT-4o Achiam et al. (2023). Besides, for a convincing comparison, an enhanced ReKep+ is introduced, where we inject an oracle-level initial grasp label to mitigate the impact of noisy perception. We also adapt two one-shot single-arm manipulation methods Mechanisms Mao et al. (2023) and MAGIC Liu et al. (2025b) for our dual-arm tasks.
@@ 4.1 Effective and Robust Bimanual Manipulation with VLBiMan
Dynamic
Inter-
ference
Manipulation
Method
@@
Average
Success
Rate
@@
Average
Success
Rate
  No Mechanisms 09/20 07/20 05/20 04/20 06/20 03/20 28.3% 05/20 04/20 01/20 01/20 03/20 01/20 12.5%
MAGIC 13/20 12/20 08/20 08/20 08/20 06/20 45.8% 09/20 08/20 04/20 04/20 05/20 04/20 28.3%
Robot-ABC 11/20 08/20 07/20 06/20 07/20 05/20 36.7% 09/20 07/20 02/20 02/20 06/20 03/20 24.2%
ReKep 11/20 09/20 08/20 10/20 08/20 06/20 43.3% 10/20 07/20 04/20 05/20 06/20 05/20 30.8%
ReKep+ 15/20 15/20 10/20 14/20 15/20 09/20 65.0% 12/20 10/20 07/20 08/20 09/20 06/20 43.3%
VLBiMan 19/20 18/20 16/20 17/20 17/20 15/20 85.0% 19/20 17/20 14/20 14/20 16/20 14/20 78.3%
  Yes Mechanisms 04/20 04/20 02/20 02/20 03/20 01/20 13.3% 02/20 01/20 00/20 00/20 01/20 00/20 3.3%
MAGIC 07/20 07/20 04/20 03/20 05/20 03/20 24.2% 04/20 03/20 02/20 01/20 03/20 01/20 11.7%
Robot-ABC 06/20 05/20 03/20 02/20 04/20 02/20 18.3% 04/20 02/20 00/20 00/20 02/20 00/20 6.7%
ReKep 08/20 05/20 05/20 03/20 04/20 03/20 23.3% 07/20 03/20 02/20 01/20 03/20 01/20 14.2%
ReKep+ 10/20 08/20 07/20 06/20 07/20 07/20 37.5% 08/20 06/20 04/20 03/20 05/20 04/20 25.0%
VLBiMan 15/20 13/20 15/20 15/20 14/20 12/20 70.0% 15/20 11/20 12/20 11/20 12/20 10/20 59.2%
@@
Dynamic
Inter-
ference
Manipulation
Method
@@
Average
Success
Rate
@@
Average
Success
Rate
  No Mechanisms 04/20 03/20 01/20 01/20 11.3% 01/20 01/20 00/20 00/20 2.5%
MAGIC 07/20 06/20 03/20 03/20 23.8% 04/20 03/20 01/20 01/20 11.3%
Robot-ABC 05/20 05/20 02/20 03/20 18.8% 03/20 02/20 00/20 01/20 7.5%
ReKep 06/20 06/20 04/20 03/20 23.8% 04/20 03/20 01/20 00/20 10.0%
ReKep+ 09/20 08/20 06/20 05/20 35.0% 06/20 05/20 03/20 02/20 20.0%
VLBiMan 12/20 12/20 10/20 08/20 52.5% 10/20 09/20 08/20 06/20 41.3%
  Yes Mechanisms 01/20 01/20 00/20 00/20 2.5% 00/20 00/20 00/20 00/20 0.0%
MAGIC 03/20 02/20 03/20 01/20 11.3% 02/20 01/20 01/20 00/20 5.0%
Robot-ABC 02/20 01/20 02/20 02/20 8.8% 01/20 00/20 01/20 00/20 2.5%
ReKep 05/20 04/20 03/20 02/20 17.5% 03/20 02/20 00/20 00/20 6.3%
ReKep+ 07/20 06/20 04/20 03/20 25.0% 05/20 03/20 01/20 01/20 12.5%
VLBiMan 10/20 09/20 07/20 05/20 38.8% 07/20 07/20 04/20 02/20 25.0%
@@ 4.3 System Performance Ablation and Analysis
  VLMs type
initial grasp
alignment
IK
refinement
collision
avoidance
Avg.
SR
@@ 5 Conclusion and Limitation
This supplementary part provides detailed clarifications and additional insights to support the main paper. In Sec. A (Discussion on Bimanual Manipulation Tasks), we present the motivation behind the design of ten bimanual manipulation tasks, including an overview of the dual-arm robotic platform, a task-by-task breakdown, and the process of collecting one-shot demonstrations for each task. In Sec. B (More Implementation Details of VLBiMan), we elaborate on the implementation details of the proposed VLBiMan framework, such as the object principal axis extraction algorithm based on image moments, and procedure for robust dual-arm execution under external dynamic disturbances. In Sec. C (More Exploration on VLBiMan Advantages and Limitations) explores further strengths of VLBiMan, including its robustness to lighting variations and its modular structure, which allows for synchronous dual-arm sub-skills to improve manipulation efficiency. We also provide additional experimental results and analyses, such as evaluating the impact of varying levels of external interference on task success rates, revealing the ease with which the system can be transferred across embodiments, as well as discussing some interesting empirical findings. Finally, Sec. D is the statement on the use of LLMs.
@@ A.3 One-shot Human Demonstration
Appendix B More Implementation Details of VLBiMan
B.1 Image Moments based Orientation Estimation
In the Vision-Language Anchored Adaptation pipeline of VLBiMan, our method requires extracting the principal axis and determining the orientation of direction-sensitive objects. This includes the marker pen in the plugpen and inserting tasks, the spoon in the reorient and tool-use spoon tasks, as well as the horizontally placed bottle in the reorient+unscrew task. As shown in Alg. 1, we adopt an object principal axis extraction algorithm based on image moments theory Chaumette (2004); Kotoulas & Andreadis (2007). Since this algorithm relies primarily on the 2D segmentation mask of object and does not require any deep networks, its computational overhead is minimal and can be considered negligible in practice.
In general, the proposed algorithm estimates the orientation angle of an object from its 2D binary mask through a hierarchical analysis of geometric properties. First, the object’s contour is extracted, and its centroid is computed using image moments. A covariance matrix derived from second-order central moments is then diagonalized to identify the principal axis direction via eigen decomposition. To resolve directional ambiguity inherent to eigenvectors, contour points are projected onto the principal axis to locate two extreme endpoints. The front endpoint is determined by comparing local perpendicular widths around these endpoints, leveraging the observation that structural asymmetry often manifests as width variation. Finally, the principal axis direction is reoriented to align with the front endpoint, and the orientation angle is calculated as the arctangent of the adjusted axis vector, ensuring a continuous 0–360 representation. This approach robustly handles directional ambiguity while maintaining computational efficiency through moment-based feature extraction.
B.2 Dynamic Interference Robust VLBiMan
Specifically, for each object to be manipulated, VLBiMan first performs continuous 2D instance segmentation and tracks the object across frames using a lightweight vision pipeline. Let $\mathbf{M}_{t}$ denote the segmentation mask at time step $t$, and let the corresponding object pose estimation function be $\mathcal{F}_{\texttt{pose}}\rightarrow(\mathbf{p}_{t},\theta_{t})$, where $\mathbf{p}_{t}$ is the 2D position and $\theta_{t}$ is the principal axis orientation obtained via image moments (refer Alg. 1). This estimation is continuously updated and serves as input to the grasp planning module. A grasping attempt is initiated only when the variance of $\{\mathbf{p}_{t-k},\cdots,\mathbf{p}_{t}\}$ and $\{\theta_{t-k},\cdots,\theta_{t}\}$ over a short sliding window falls below a pre-defined threshold $\epsilon$ (e.g., absolute moving distance less than 10mm), indicating that the object has stabilized. This implicitly filters out moments of dynamic perturbation. Once the object is deemed stable, the robot executes the corresponding grasp action $\mathcal{G}(\mathbf{p}_{t},\theta_{t})$, where $\mathcal{G}(\cdot)$ denotes a grasp generation function conditioned on both position and orientation. If the grasp fails (e.g., the object slips or moves significantly post-action), the system returns to the observation loop and restarts the stabilization-checking process. This mechanism ensures that the object’s interaction policy is dynamically robust, without requiring hard-coded assumptions on when or how disturbances may occur.
Dynamic
Interference
Uneven
Lighting
new placements + same objects new placements + novel instances
Manipulation
Method
@@ B.2 Dynamic Interference Robust VLBiMan
Average
Success
Rate

### v3

@@ 2 Related Works
Generalizable Representations for Manipulation. Traditional robotic manipulation often relied on structured representations built upon strong priors Kaelbling and Lozano-Pérez (2013); Dantam et al. (2018); Migimatsu and Bohg (2020); Tyree et al. (2022), such as object geometry or rigid-body assumptions, typically via estimating 6D poses or manually specifying grasp configurations. They are hard to scale in unstructured environments. With the rise of data-driven techniques, more flexible representations have emerged, including keypoints Papagiannis et al. (2024); Gao et al. (2024); Wen et al. (2024b); Grannen et al. (2021), affordances Ju et al. (2024); Nasiriany et al. (2024); Zhao et al. (2023b), dynamic flow fields Colomé and Torras (2018); Weng et al. (2022), and invariant object-centric correspondences Ko et al. (2024); Zhang and Boularias (2024); Zhang et al. (2023). Some works further leverage human demonstrations to retarget 3D hand trajectories to robots Chen et al. (2024a); Li et al. (2024); Kerr et al. (2024); Chen et al. (2024b). However, these approaches often depend on private datasets, retraining, or complex retargeting pipelines, limiting scalability. In contrast, our method essentially anchors adaptation to object representive points without retraining, achieving greater efficiency and generality.
@@
Efficient Bimanual Robotic Manipulation. Recent advances in bimanual manipulation have showcased the power of large Vision-Language-Action (VLA) models Black et al. (2024); Liu et al. (2025a); Pertsch et al. (2025) trained on extensive teleoperated demonstrations Fang et al. (2024a); Khazatsky et al. (2024); O’Neill et al. (2024); Bu et al. (2025). However, these approaches are highly suspected of lacking efficiency, as scaling to unseen objects or tasks often requires re-collecting and retraining. Alternative efforts explore leveraging large-scale Internet Ponimatkin et al. (2025); Ye et al. (2025); Bharadhwaj et al. (2024) or egocentric human-hand videos Zhan et al. (2024); Liu et al. (2024b); Grauman et al. (2024); Zhao et al. (2025); Kareer et al. (2024), yet the embodiment gap between human and robot limits direct usability. Some methods improve sample efficiency by learning visuomotor policies Chi et al. (2023); Ze et al. (2024) from a small set of real-world robot data, but their generalization remains limited. While one-shot imitation learning Wen et al. (2022); Bahety et al. (2024); Zhou et al. (2025); Wang and Johns (2025); Mao et al. (2023); Liu et al. (2025b); Biza et al. (2023) reduces data demands, the high-dimensional action space and coordination complexity of bimanual control hinder learning efficiency. In contrast, VLBiMan achieves efficient adaptation from a single bimanual demonstration by leveraging VLMs to handle novel variations, while reusing decomposed task-invariant atomic skills. These lead to both data and computational efficiency.
@@
Large Foundation Models for Robotics. Integrating LLMs and VLMs into robotics is a prominent trend to enable generalizable agents Ma et al. (2025); Huang et al. (2025); Fang et al. (2025); Feng et al. (2025). LLMs are utilized for high-level task understanding and planning, such as decomposing instructions into executable subtasks or generating scripts Liang et al. (2023); Singh et al. (2023); Szot et al. (2024); Huang et al. (2024a). Meanwhile, VLMs facilitate visually grounded perception through semantic prompts, enabling object-level detection and segmentation. For fine-grained tasks, Visual Foundation Models (VFMs) Oquab et al. (2024); Ravi et al. (2025) are further employed to find keypoints Papagiannis et al. (2024); Gao et al. (2024); Wen et al. (2024b) or dense correspondences Ko et al. (2024); Zhang and Boularias (2024). Recent efforts like ReKep Huang et al. (2024b), MOKA Fang et al. (2024b), RobotPoint Yuan et al. (2024a), and RAM Kuang et al. (2024) combine LLMs and VLMs into modular pipelines that follow the perceive-understand-plan-act paradigm to achieve zero-shot generalization. These approaches often rely on engineered prompts and ambiguous intermediate representations (e.g., region of interest or keypoint clusters) requiring additional post-processing. In contrast, VLBiMan avoids LLM-based instruction parsing and task decomposition, which are brittle and labor-intensive. Instead, we build on one-shot demonstrations with precise action labels, using VLMs to extract semantically grounded action structures that are adaptively composed and reused, enabling efficient and scalable bimanual manipulation.
@@ 3.2 Vision-Language Anchored Adaptation
VFM-Based Geometric Feasibility. To adapt grasping or alignment poses, we introduce a three-step process. (1) Firstly, we compute relative position transformation $\Delta\mathbf{T}$ between new object placement and reference demonstration via task-specific representative points (e.g. the 2D mask centroid or a task-specific contact point on the table-facing boundary). Examples of two kinds of representative points can be found in Fig. 3. Let $\mathbf{p}^{\texttt{demo}}$, $\mathbf{p}^{\texttt{new}}$ denote the representative 3D positions back-projected from 2D points via stereo and calibrated camera intrinsics. The relative position shift is $\Delta\mathbf{x}=\mathbf{p}^{\texttt{new}}-\mathbf{p}^{\texttt{demo}}$. (2) To account for orientation-sensitive objects (such as pen, spoon and lying down bottle), we compute the principal axis from second-order image moments Chaumette (2004); Kotoulas and Andreadis (2007) of the 2D mask and derive relative rotation $\Delta\theta=\angle(\mathbf{v}^{\texttt{new}},\mathbf{v}^{\texttt{demo}})$ via angular deviation. The final adapted grasp pose $\widetilde{\mathbf{T}}$ is obtained by applying $(\Delta\mathbf{x},\Delta\theta)$ to the original grasp pose in robot coordinates via calibrated hand-eye transformation. (3) For category-level variation, we measure shape-induced feasibility change through height and width differences. For example, the $z$-extent of the object point cloud $\mathcal{P}_{k}^{\texttt{3D}}$ yields $\Delta h_{k}=\max_{z}(\mathcal{P}_{k}^{\texttt{3D}})-\min_{z}(\mathcal{P}_{k}^{\texttt{3D}})$. This is used to adjust vertical placement motions or inter-arm distances for tools or containers.
@@ 4 Experiments
Baselines and Metric: For each task setting, we conduct 25 trials, where objects are randomly located or replaced, and the success rate will be reported. For baselines, we compare to Robot-ABC Ju et al. (2024) based on keypoint affordance prediction with using AnyGrasp Fang et al. (2023) for initial grasping (After which, the remaining trajectory is obtained by trivial modules combination), as well as ReKep Huang et al. (2024b) based on VFMs (SAM Kirillov et al. (2023) and DINOv2 Oquab et al. (2024)) and GPT-4o Achiam et al. (2023). Besides, for a convincing comparison, an enhanced ReKep+ is introduced, where we inject an oracle-level initial grasp label to mitigate the impact of noisy perception. We also adapt two one-shot single-arm manipulation methods Mechanisms Mao et al. (2023) and MAGIC Liu et al. (2025b) for our dual-arm tasks.
@@ 4.1 Effective and Robust Bimanual Manipulation with VLBiMan
Dynamic Inter- ference Manipulation Method
@@
Average Success Rate
@@
Average Success Rate
  No Mechanisms 11/25 09/25 05/25 05/25 07/25 03/25 26.7% 06/25 05/25 02/25 01/25 04/25 01/25 12.7%
MAGIC 16/25 15/25 10/25 10/25 09/25 07/25 44.7% 11/25 10/25 05/25 05/25 06/25 04/25 27.3%
Robot-ABC 14/25 10/25 09/25 07/25 08/25 06/25 36.0% 11/25 09/25 03/25 02/25 07/25 04/25 24.0%
ReKep 14/25 11/25 10/25 12/25 10/25 08/25 43.3% 12/25 08/25 05/25 06/25 07/25 06/25 29.3%
ReKep+ 19/25 18/25 13/25 17/25 17/25 11/25 63.3% 15/25 12/25 09/25 10/25 11/25 07/25 42.7%
VLBiMan 25/25 23/25 20/25 21/25 20/25 19/25 85.3% 24/25 21/25 18/25 17/25 20/25 17/25 78.0%
  Yes Mechanisms 05/25 05/25 03/25 02/25 04/25 01/25 13.3% 03/25 01/25 00/25 00/25 02/25 00/25 4.0%
MAGIC 09/25 09/25 05/25 04/25 06/25 04/25 24.7% 05/25 04/25 03/25 01/25 04/25 01/25 12.0%
Robot-ABC 07/25 06/25 04/25 03/25 05/25 02/25 18.0% 05/25 03/25 00/25 00/25 03/25 00/25 7.3%
ReKep 10/25 06/25 06/25 04/25 05/25 03/25 22.7% 09/25 04/25 03/25 01/25 04/25 02/25 15.3%
ReKep+ 12/25 10/25 09/25 08/25 09/25 09/25 38.0% 10/25 08/25 05/25 04/25 06/25 05/25 25.3%
VLBiMan 19/25 16/25 19/25 18/25 17/25 15/25 69.3% 18/25 14/25 15/25 14/25 15/25 13/25 59.3%
@@
Dynamic Inter- ference Manipulation Method
@@
Average Success Rate
@@
Average Success Rate
  No Mechanisms 05/25 04/25 02/25 01/25 12.0% 01/25 02/25 00/25 00/25 3.0%
MAGIC 09/25 08/25 04/25 03/25 24.0% 05/25 04/25 01/25 01/25 11.0%
Robot-ABC 06/25 06/25 03/25 03/25 18.0% 04/25 02/25 00/25 01/25 7.0%
ReKep 07/25 08/25 05/25 03/25 23.0% 05/25 04/25 01/25 00/25 10.0%
ReKep+ 11/25 10/25 07/25 06/25 34.0% 07/25 06/25 04/25 02/25 19.0%
VLBiMan 15/25 15/25 12/25 10/25 52.0% 12/25 11/25 10/25 08/25 41.0%
  Yes Mechanisms 01/25 02/25 00/25 00/25 3.0% 00/25 00/25 00/25 00/25 0.0%
MAGIC 04/25 03/25 04/25 01/25 12.0% 02/25 02/25 01/25 00/25 5.0%
Robot-ABC 02/25 02/25 02/25 02/25 9.0% 01/25 00/25 01/25 00/25 2.0%
ReKep 06/25 05/25 03/25 02/25 16.0% 03/25 03/25 00/25 00/25 6.0%
ReKep+ 08/25 08/25 05/25 03/25 24.0% 06/25 04/25 01/25 01/25 12.0%
VLBiMan 12/25 11/25 09/25 06/25 38.0% 08/25 09/25 05/25 02/25 24.0%
@@ 4.3 System Performance Ablation and Analysis
  VLMs type initial grasp alignment IK refinement collision avoidance Avg. SR
@@ 5 Conclusion and Limitation
Acknowledgments
This work was supported by the Guangdong Provincial Key Field R&D Program (Project No. 20240104), and was also funded by the Shenzhen Science and Technology Major Project (No. 202402002 and ZDCT20250901113000001).
@@ Acknowledgments
This supplementary part provides detailed clarifications and additional insights to support the main paper. In Sec. A (Discussion on Bimanual Manipulation Tasks), we present the motivation behind the design of ten bimanual manipulation tasks, including an overview of the dual-arm robotic platform, a task-by-task breakdown, and the process of collecting one-shot demonstrations for each task. In Sec. B (More Implementation Details of VLBiMan), we elaborate on the implementation details of the proposed VLBiMan framework, such as the object principal axis extraction algorithm based on image moments, and procedure for robust dual-arm execution under external dynamic disturbances. In Sec. C (More Exploration on VLBiMan Advantages and Limitations), we explore further strengths of VLBiMan, including its robustness to lighting variations and its modular structure, which allows for synchronous dual-arm sub-skills to improve manipulation efficiency. We also provide additional experimental results and analyses, such as evaluating the impact of varying levels of external interference on task success rates, revealing the ease with which the system can be transferred across embodiments, as well as discussing some interesting empirical findings. This section also includes additional analyses introduced in four new subsections: (1) a detailed discussion of the human-in-the-loop refinement process used during primitive segmentation, clarifying its role and negligible burden; (2) an investigation into the robustness of using simple object representing points—such as mask centroids or front-edge contacts—for cross-object generalization; (3) evaluations of VLBiMan under cluttered scenes to assess stability in more realistic environments; and (4) ablation studies on pre-grasp interpolation density, highlighting its effect on collision avoidance and resilience to external disturbances. Finally, Sec. D is the statement on the use of LLMs.
@@ A.3 One-shot Human Demonstration
Table 4: Statistical details regarding the ten bimanual manipulation tasks defined in this study. They contain the names of the target objects involved (including their placement states), the representation points used to indicate the positions of the target objects (where MC represents the object’s 2D mask centroid, and CP represents the contact point between the object bottom and the table top. Please refer to Fig. 3 in the main text for visualization), and whether the object’s orientation needs to be estimated during the manipulation process.
  Task Name
@@
reorient +unscrew
unscrew +pouring
tool-use scoop
tool-use funnel
  Object Name marker body marker cap marker pen cup standing bottle standing bottle mug cup pump bottle spoon or shovel lying bottle standing bottle mug spoon big bowl small bowl funnel standing bottle mug
Representive Point Type MC MC MC CP CP CP CP CP CP MC MC CP CP MC CP CP CP CP CP
Orientation Estimation?
Appendix B More Implementation Details of VLBiMan
B.1 Image Moments based Orientation Estimation
In the Vision-Language Anchored Adaptation pipeline of VLBiMan, our method requires extracting the principal axis and determining the orientation of direction-sensitive objects. This includes the marker pen in the plugpen and inserting tasks, the spoon in the reorient and tool-use spoon tasks, as well as the horizontally placed bottle in the reorient+unscrew task. As shown in Algorithm 1, we adopt an object principal axis extraction algorithm based on image moments theory Chaumette (2004); Kotoulas and Andreadis (2007). Since this algorithm relies primarily on the 2D segmentation mask of object and does not require any deep networks, its computational overhead is minimal and can be considered negligible in practice.
In general, the proposed algorithm estimates the orientation angle of an object from its 2D binary mask through a hierarchical analysis of geometric properties. First, the object’s contour is extracted, and its centroid is computed using image moments. A covariance matrix derived from second-order central moments is then diagonalized to identify the principal axis direction via eigen decomposition. To resolve directional ambiguity inherent to eigenvectors, contour points are projected onto the principal axis to locate two extreme endpoints. The front endpoint is determined by comparing local perpendicular widths around these endpoints, leveraging the observation that structural asymmetry often manifests as width variation. Finally, the principal axis direction is reoriented to align with the front endpoint, and the orientation angle is calculated as the arctangent of the adjusted axis vector, ensuring a continuous 0–360 representation. This approach robustly handles directional ambiguity while maintaining computational efficiency through moment-based feature extraction. For more details on which objects in which tasks require the orientation estimation algorithm, please refer to Tab. 4 (see the last row).
Why Not Use Off-the-Shelf 6D Pose Estimation? While one may consider leveraging off-the-shelf 6D pose estimators Lin et al. (2024); Wen et al. (2024a) for object orientation extraction, we found that such solutions are unnecessary, unstable across objects, and incompatible with VLBiMan’s cross-object generalization objective. Classical and learning-based 6D pose estimators generally require either (1) object-specific CAD models, (2) textured templates, or (3) category-level canonicalization priors. These assumptions are difficult to satisfy in our setup, where VLBiMan must generalize to unseen and shape-diverse everyday objects without additional training or model registration. Moreover, 6D estimators often degrade when objects lack distinctive geometry or texture—precisely the case for many household items used in our tasks (e.g., plain spoons, cylindrical pens). In contrast, our moment-based orientation estimation (Algorithm 1) only depends on the 2D segmentation mask produced by a VLM-powered perception module, eliminating the need for any object-specific shape information. This makes the approach far more robust to appearance variations and naturally compatible with VLBiMan’s object-centric anchoring framework. Additionally, we observed in our experiments that 6D pose estimators frequently output unstable yaw angles under partial occlusion or when only a single RGB camera view is available, while the moment-based method remains consistent, lightweight, and easy to deploy in real-world bimanual settings.
Finally, this simplified 2D-mask–driven orientation strategy is fully aligned with VLBiMan’s design principle—to avoid heavy perception modules that compromise generalization—and it has proven sufficient for all direction-sensitive tasks, including plugpen, inserting, reorient, tool-use spoon, and reorient+unscrew. Hence, the choice to avoid 6D pose estimation is both practical and necessary: our goal is not to recover a full metric pose, but to obtain a stable, VLM-compatible orientation anchor that enables one-shot bimanual manipulation without retraining or object modeling.
B.2 Dynamic Interference Robust VLBiMan
Specifically, for each object to be manipulated, VLBiMan first performs continuous 2D instance segmentation and tracks the object across frames using a lightweight vision pipeline. Let $\mathbf{M}_{t}$ denote the segmentation mask at time step $t$, and let the corresponding object pose estimation function be $\mathcal{F}_{\texttt{pose}}\rightarrow(\mathbf{p}_{t},\theta_{t})$, where $\mathbf{p}_{t}$ is the 2D position and $\theta_{t}$ is the principal axis orientation obtained via image moments (refer Algorithm 1). This estimation is continuously updated and serves as input to the grasp planning module. A grasping attempt is initiated only when the variance of $\{\mathbf{p}_{t-k},\cdots,\mathbf{p}_{t}\}$ and $\{\theta_{t-k},\cdots,\theta_{t}\}$ over a short sliding window falls below a pre-defined threshold $\epsilon$ (e.g., absolute moving distance less than 10mm), indicating that the object has stabilized. This implicitly filters out moments of dynamic perturbation. Once the object is deemed stable, the robot executes the corresponding grasp action $\mathcal{G}(\mathbf{p}_{t},\theta_{t})$, where $\mathcal{G}(\cdot)$ denotes a grasp generation function conditioned on both position and orientation. If the grasp fails (e.g., the object slips or moves significantly post-action), the system returns to the observation loop and restarts the stabilization-checking process. This mechanism ensures that the object’s interaction policy is dynamically robust, without requiring hard-coded assumptions on when or how disturbances may occur.
  Dynamic Interference + Uneven Lighting new placements + same objects new placements + novel instances
Manipulation Method